# London, UK Dual Branch VideoMamba for Violence Detection

### v1

@@ 3.3 Final Fusion Block
$\mathbf{Z}=\operatorname{concat}(CLS_{fin}^{1},CLS_{fin}^{2}),$ (8)
Finally, the target class Pr is obtained by passing Z through a fully connected projection layer.
@@ 4.1 Datasets
The most challenging datasets so far in the real-world violence detection domain are the Real-World Fighting (RWF-2000) dataset [Cheng et al.(2021)Cheng, Cai, and Li], Real Life Violence Situations (RLVS) dataset [Soliman et al.(2019)Soliman, Kamal, Nashed, Mostafa, Chawky, and Khattab], and VioPeru dataset [Huillcen Baca et al.(2024)Huillcen Baca, Palomino Valdivia, and Gutierrez Caceres] that contain video footage of fighting in real life scenarios. Out of these three datasets, while the RWF-2000 and VioPeru datasets contain exclusive surveillance footage, RLVS contains violence action which includes surveillance footage as well. All three datasets contain videos cropped at 5s and depict real-world scenarios. We considered other violence-related datasets, however, we observed these have many clips which are not strongly labeled or acquired from staged violence. Therefore, for the rest of the paper, we focus on these three data sets with the intention of assessing our model to reflect real-world violence scenarios.
4.1.1 Combined Dataset of RWF-2000, RLVS, and VioPeru
In order to create a more comprehensive and challenging benchmark for violence detection, we combine the three datasets: RWF-2000, RLVS, and VioPeru. Although each dataset has its own predefined train-test split, simply merging them without ensuring strict partition integrity can lead to potential data leakage, where identical videos appear in both the training and testing sets, since RWF-2000 and RLVS were both sourced from YouTube  [Cheng et al.(2021)Cheng, Cai, and Li, Soliman et al.(2019)Soliman, Kamal, Nashed, Mostafa, Chawky, and Khattab]. To address this concern, we embed all videos from the training and testing splits of each dataset using VideoMAE  [Tong et al.(2022)Tong, Song, Wang, and Wang] embeddings. Next, we compute the cosine similarity between every pair of videos (across training and testing splits) to detect duplicates. All the pairs which yielded 78% or higher cosine similarity were manually inspected.
Final Combined Dataset Statistics: Duplicates in the training sets were not investigated because they naturally represent the frequency of common real-world situations, ensuring the model learns an accurate distribution. Also, since regularization techniques are used in our model training, we can safely assume that biased learning from duplicate data is mitigated. After eliminating the identified duplicates from the RLVS test split, the statistics of final combined dataset is reported in Table 1.
@@ 4.1.1 Combined Dataset of RWF-2000, RLVS, and VioPeru
By merging these three datasets while enforcing a strict data leakage check, we obtain a unified benchmark that provides diverse scenarios of violent and non-violent events, captured in varying resolutions and environmental conditions. This amalgamated dataset thus serves as a more robust testbed for evaluating the generalizability and performance of violence detection models.
Implementation Details: Our algorithm was implemented with the specifications of the VideoMamba-Middle architecture [Li et al.(2024a)Li, Li, Wang, He, Wang, Wang, and Qiao] with 32 layers where the hidden dimension $d$ was 576 across the layers, for each branch in order to facilitate loading pre-trained weights of the vanilla VideoMamba trained with Kinetics-400 dataset [Kay et al.(2017)Kay, Carreira, Simonyan, Zhang, Hillier, Vijayanarasimhan, Viola, Green, Back, Natsev, et al.]. Further details can be found in the Section 6.4 of supplementary material.
@@ 4.2 Results
Arch. Model Pretraining Input Size Params (M) FLOPS (G) Top-1 Acc. (%) F1-Viol.(%) F1-Non-Viol.(%)
CNN SlowFast - 64$\times$2242 60 234 76.37 74.56 77.95
Trans. VideoSwin-B K-400 64$\times$2242 88 281.6 80.47 79.15 81.63
CNN + Trans. Uniformer-V2 K-400 64$\times$2242 354 6108 92.51 92.87 92.12
SSM VideoMamba K-400 64$\times$2242 74 806 94.39 94.29 94.48
CNN + Trans. CUE-Net K-400 64$\times$2242 354 5826 95.91 95.95 95.86
SSM Dual Branch VideoMamba K-400 64$\times$2242 154.3 1612 96.37 96.33 96.42
Table 2: Comparative results among various architectures for the newly combined dataset. Each model is characterized by its backbone type, the pretraining dataset, the input resolution, the number of parameters, the FLOPS, the Top-1 accuracy (%), and the F1-scores (%) for violent and non-violent classes.
In this section we perform an in-depth analysis comparing our Dual Branch VideoMamba architecture with other leading architectures using the combined dataset. Following the standard practice used for the three combined datasets, we also use classification accuracy as the main metric to evaluate the performance. The results in Table 2 show a clear progression in performance across the listed architectures. While the SlowFast model [Feichtenhofer et al.(2019)Feichtenhofer, Fan, Malik, and He], VideoSwin Transformer [Liu et al.(2022)Liu, Ning, Cao, Wei, Zhang, Lin, and Hu] and UniFormer-V2 [Li et al.(2023b)Li, Wang, He, Li, Wang, Wang, and Qiao] demonstrate respectable accuracies, they are outperformed by VideoMamba [Li et al.(2024a)Li, Li, Wang, He, Wang, Wang, and Qiao] and CUE-Net [Senadeera et al.(2024)Senadeera, Yang, Kollias, and Slabaugh]. Notably, our Dual Branch VideoMamba model attains the highest scores, with a test accuracy of 96.37% and corresponding F1-scores of 96.33% and 96.42% for violence and non-violence, respectively. Furthermore, these top accuracies are achieved with a model size of 154 M parameters and a FLOPS count of 1612 GFLOPS, compared to the CUE-Net model which has 354 M parameters and 5826 GFLOPS. This represents more than a 50% reduction in both the number of parameters and the FLOPS count. Such improvements highlight the efficiency of our dual-branch SSM design over traditional CNN+Transformer architectures. Performance of our architecure on each individual dataset is detailed in Section 6.3 of supplementary material.
@@
Figure 4: Visual Analysis of Class Activation Maps (CAMs). (a) CAMs for two violent-labelled videos correctly classified as violent. (b) CAMs for two non-violent videos misclassified as violent focusing on a forceful object removal and a collision.
Visual Analysis of Results: We use Grad-CAM [Selvaraju et al.(2017)Selvaraju, Cogswell, Das, Vedantam, Parikh, and Batra] to visualize class activation maps (CAMs). As SSM models motion through internal state evolution rather than spatial saliency, temporal interpretability remains a key challenge. So we analyze CAMs from the spatial branch to partially address this. The spatial-first branch effectively highlights object-level cues via spatial gradients in Mamba blocks. In Fig. 4(a), the CAMs for two violent-labeled videos that were correctly classified as violent, show that the model accurately focuses on regions where human interactions which are indicative of violence occur, even when these interactions are partially occluded. In contrast, Fig. 4(b) shows CAMs for two non-violent videos that the model mistakenly classified as violent due to a forceful object removal and an accidental collision. These observations indicate that, while the model has effectively learned to detect violent cues, it may occasionally overemphasize certain visual signals too.
@@ 4.3 Ablation Study
We performed a series of ablation studies to assess the efficacy of the components of our architecture. Table 4 compares model performance with and without the cropping module. We observe that the architecture benefit from the cropping mechanism, as it focuses the network on the regions where people are most likely to appear. In Table 4, we investigate the optimal configuration for lateral connections between the two branches. Continuous lateral connections (i.e., applying them at every block) yield the highest accuracy of 96.37% where this finding highlights the importance of gradually fusing information from both branches continuously throughout the network rather than restricting fusion to early or late stages. In Table 5, we compare 3 fusion mechanisms for lateral connections: gated, additive, and concatenation-based approaches where Gated Lateral Connections (from Branch-1 to Branch-2) outperform all other methods. These results indicate that the Gated Class Token Fusion (GCTF) mechanism is highly effective in adaptively blending complementary features between branches and also shows that the direction of information flow from spatial-first scanning to temporal-first scanning also plays a critical role in maximizing performance.
Model Cropping Module Accuracy (%)
VideoMamba $\times$ 94.39
VideoMamba $\checkmark$ 95.79
Dual Branch VideoMamba $\times$ 95.44
Dual Branch VideoMamba $\checkmark$ 96.37
Table 3: Performance comparison with and without the cropping module.
Lateral-Connection-(LC) Freq. Accuracy (%)
LCs alternatively (even layers) 95.78
LCs alternatively (odd layers) 95.78
One LC only at beginning 95.44
One LC only at end 95.67
One LC at middle 95.56
Two LCs at the end and beginning 95.91
Continuous LCs 96.37
Table 4: Performance for different configurations of lateral connections.
@@
Additive LCs 94.78
Concatenated LCs 95.91
Gated LCs (Branch-2 to Branch-1) 94.04
Gated LCs (Branch-1 to Branch-2) 96.37
Table 5: Performance comparison for different fusion mechanisms.
@@ 5 Conclusion
In this paper, we present a novel Dual Branch VideoMamba architecture with Gated Class Token Fusion (GCTF) for violence detection in surveillance based videos. Our method integrates an efficient state-space model with a dual-stream design aimed at separately capturing spatial features and temporal dynamics. By continuously fusing the class tokens from both branches using a learnable gating mechanism, our approach effectively combines complementary cues, yielding state-of-the-art performance. Extensive experiments on the combined benchmark dataset demonstrate that our model not only achieves state-of-the-art accuracies but also shows a significant reductions in model parameters and computational cost in FLOPS compared to previous state-of-the-art methods such as CUE-Net [Senadeera et al.(2024)Senadeera, Yang, Kollias, and Slabaugh]. This work highlights the promise of state-space models for scalable and real-time video violence detection in surveillance and paves the way for future research, including the integration of multi-modal data and further refinement of the fusion strategies.
@@
Acknowledgment. This work has been funded through an EPSRC DTP studentship at Queen Mary University of London in collaboration with Remark AI UK Ltd. This paper utilised Queen Mary’s Andrena HPC facility, supported by QMUL Research-IT Services [King et al.(2017)King, Butcher, and Zalewski].

### v2

@@ 3.3 Final Fusion Block
$\mathbf{Z}=\operatorname{concat}(\text{CLS}_{fin}^{1},\text{CLS}_{fin}^{2}),$ (8)
Finally, the target class $Pr$ is obtained by passing Z through a fully connected projection layer. Further details are discussed in  Sec. 7.1.2 of supplementary material
Figure 3: (a) Spatial-First Scanning (b) Temporal-First Scanning.
@@ 4.1 Datasets
Architecture Model Pretraining Input Size Params (M) FLOPS (G) Combined Dataset (%) DVD Dataset (%)
Top-1 F1-V F1-NV Top-1 F1-V F1-NV
[…]
Table 1: Comparative results among various architectures for the newly combined dataset and the DVD dataset. Each model is characterized by its backbone type (CNN, Transformer, SSM), the pretraining dataset, the input resolution, the number of parameters (in millions), the FLOPS, the Top-1 Accuracy (%), and the F1-scores (%) for Violent and Non-Violent classes.
4.1.1 Combined Dataset of RWF-2000, RLVS, SURV, and VioPeru
Until recently, the most challenging strongly labeled open-source datasets in the real-world violence detection domain are the Real-World Fighting (RWF-2000) dataset [9], Real Life Violence Situations (RLVS) dataset [36], Vision-based Fight Detection From Surveillance Cameras (SURV) dataset [2], and the VioPeru dataset [16], which contain video footage of fighting in real-life scenarios. Out of these four datasets, the RWF-2000, SURV, and VioPeru datasets contain exclusive surveillance footage. All four datasets consist of short trimmed clips (2–5 seconds), depicting real-world scenarios in contrast with other strongly labeled violence datasets such as the AUTO dataset [6], which instead relies on actors staging violent incidents under controlled conditions. Furthermore, we did not incorporate the Movie Scene Violence [32] dataset or the Hockey Fight [5] dataset, as they are purely sourced from films and sports recordings, and therefore do not adequately represent real-world violent encounters.
We also did not consider the UCF-Crime dataset [39] and the XD-Violence dataset [47] in creating this amalgamation, because their training data are weakly labeled: both datasets provide only video-level labels (Violent / Non-Violent) without precise annotations of when violent actions occur within the clips, making them unsuitable to mix with strongly labeled data. Therefore, we decided to move forward with these four datasets with the intention of assessing our model not on staged violent scenarios, but on strongly labeled real-world violent videos. Although each dataset has its own predefined train-test split, simply merging them without ensuring strict partition integrity can lead to potential data leakage, where identical videos appear in both the training and testing sets, since RWF-2000 and RLVS were both sourced from YouTube [9, 36]. To address this concern, we take the following precautions:
Cosine Similarity Check: We embed all videos from the training and testing splits of each dataset using VideoMAE [42] embeddings. Next, we compute the cosine similarity between every pair of videos (across training and testing splits) to detect duplicates. A high cosine similarity score (close to 1) would indicate that two videos are effectively identical or extremely similar in content. All the pairs which yielded 75% or higher cosine similarity were manually inspected.
Final Combined Dataset Statistics: Duplicates in the training sets were not investigated because they naturally represent the frequency of common real-world situations, ensuring the model learns an accurate distribution. Also, since regularization techniques are used in our model training, we can safely assume that biased learning from duplicate data is mitigated. After eliminating the identified duplicate from the RLVS test split, the statistics of final combined dataset is reported in Tab. 2.
@@ 4.1.1 Combined Dataset of RWF-2000, RLVS, SURV, and VioPeru
By merging these four datasets while enforcing a strict data leakage check, we obtain a unified benchmark that provides diverse scenarios of violent and non-violent events, captured in varying resolutions and environmental conditions. This amalgamated dataset thus serves as a more robust testbed for evaluating the generalizability and performance of violence detection models.
4.1.2 DVD Dataset
A novel dataset named DVD [22] has been introduced recently in the space of video violence detection. It is a large-scale dataset comprising 344 videos and 2.7M frames with frame-level annotations for violence detection. DVD dataset is designed to capture diverse environments, varying lighting conditions, multiple camera sources, and complex social interactions [22]. It is specifically designed to reflect the complexities of real-world violent events.
Following the official frame-level annotations of DVD [22], we segmented the dataset into continuous clips, where consecutive violent frames were grouped into a single violent clip, and consecutive non-violent frames were grouped into a non-violent clip. This process resulted in a total of 2,648 clips, comprising 1,099 violent clips ($\approx 9.5$ hours) and 1,549 non-violent clips ($\approx 15$ hours), spanning approximately 24.6 hours of footage. To avoid data leakage, clips originating from the same source video were strictly assigned either to the training set or to the testing set, but never to both. The final approximately 80%/20% train-test split preserved the overall balance of violent and non-violent content, with 1,987 clips (820 violent, 1,167 non-violent) in the training set and 661 clips (279 violent, 382 non-violent) in the test set, as summarized in Tab. 3.
Violent Non-Violent Total
Training Set 820 1167 1987
Testing Set 279 382 661
Total 1099 1549 2648
Table 3: Training and Testing split statistics of the DVD dataset, showing the number of violent and non-violent clips.
4.2 Implementation Details
Our algorithm was implemented with the specifications of the VideoMamba-Middle (M) architecture [26] with 32 layers where the hidden dimension $d$ was 576 across the layers, for each branch in order to facilitate loading pre-trained weights of Vanilla VideoMamba trained with Kinetics-400 dataset [20]. Further details can be found in the Sec. 6 of supplementary materials.
@@ 4.3 Results
Figure 4: Visual Analysis of Class Activation Maps (CAMs). (a) CAMs for two violent-labeled videos correctly classified as violent, illustrating that the model accurately focuses on regions of human interaction—even under occlusion and near frame edges. (b) CAMs for two non-violent videos misclassified as violent: in one case, a forceful object removal is highlighted, and in the other, a collision involving a woman and a group is emphasized.
In this section, we perform an in-depth analysis comparing our Dual Branch VideoMamba architecture with other leading architectures using the combined dataset and the DVD dataset. Following the standard practice adopted in prior work, we employ classification accuracy together with class-wise F1 scores as our primary evaluation metrics. In particular, previous work [35, 23, 9, 16] have primarily reported results using accuracy on these datasets. Therefore, for consistency and comparability, we adhere to the same evaluation protocol where we additionally include class-wise F1 scores to better capture the per-class performance in this binary setting.
The results in Tab. 1 shows a clear progression in performance across the listed architectures for the combined dataset. While the SlowFast model [11], VideoSwin Transformer [28] and UniFormer-V2 [25] demonstrate respectable accuracies, they are outperformed by the two variants of VideoMamba [26] and CUE-Net [35], each achieving above 92% and 94% test accuracies respectively for the combined dataset. Notably, our Dual Branch VideoMamba model attains the highest scores for the combined dataset, with a test accuracy of 95.85% and corresponding F1-scores of 95.89% and 95.81% for violent and non-violent classes, respectively.
The results in Tab. 1 for the DVD dataset highlight its increased difficulty, stemming from the inherent class imbalance (279 violent vs. 382 non-violent clips). While SlowFast [11], VideoSwin [28], and UniFormer-V2 [25] achieve moderate accuracies, they are surpassed by VideoMamba variants [26] and CUE-Net [35], which exceed 71% Top-1 accuracy. Notably, our Dual Branch VideoMamba again delivers the strongest results, achieving 74.13% accuracy with F1-scores of 68.97% and 77.82% for violent and non-violent classes, respectively. The higher F1 score for the majority class reflects the dataset imbalance, yet the model maintains competitive performance on the minority violent class, underscoring its robustness to real-world distributions.
Furthermore, these top accuracies are achieved with a model size of 154 M parameters and a FLOPS count of 1830 GFLOPS, compared to the CUE-Net model which has 354 M parameters and 5826 GFLOPS. This represents more than a 50% reduction in both the number of parameters and the FLOPS count. When compared to the VideoMamba - Large variant, which is of a comparable scale (148M vs. our 154M parameters), our Dual Branch VideoMamba consistently performs better, achieving 95.85% vs. 92.46% Top-1 accuracy on the combined dataset and 74.13% vs. 71.10% on the DVD dataset. This makes the comparison particularly fair, as both models operate under similar capacity constraints, yet our architecture demonstrates clear performance gains. To further validate these improvements over the VideoMamba-Large model, we carried out McNemar’s two-sided exact test [31]. Both, on the combined dataset ($n_{01}=15$, $n_{10}=46$, $d=61$, $p=8.84\times 10^{-5}$) and on the DVD dataset ($n_{01}=32$, $n_{10}=52$, $d=84$, $p=0.0375$), the differences were found to be statistically significant at $p=0.05$, confirming the robustness of our model’s advantage. Such improvements emphasize the efficiency of this dual-branch design. Performance of our architecture on each individual dataset is detailed in Sec. 7 of supplementary material.
4.4 Visual Analysis of Results
We have employed Grad-CAM [34] to generate class activation maps (CAMs) to visualize the regions that contribute most to the model’s decisions. In the spatial-first scanning branch, this approach is particularly effective because it primarily intends to extract spatial features and object-level cues via Mamba blocks which relies on spatial gradients to highlight discriminative regions. However, the temporal-first scanning branch intends to process motion dynamics over sequences of frames and relies on hidden state evolution via the Mamba mechanism rather than direct spatial activations. As a result, Grad-CAM might struggle to localize the change of motion cues over a set of frames effectively [12]. Therefore, we generate and analyze class activation maps from the spatial branch only.
In Fig. 4(a), the class activation maps (CAMs) for two violent-labeled videos that were correctly classified as violent, show that the model accurately focuses on regions where human interactions which are indicative of violence occur, even when these interactions are partially occluded or are happening near the edge of the frames. In contrast, Fig. 4(b) shows CAMs for two non-violent videos that the model mistakenly classified as violent. In the first example, the model highlights the area where a rugby ball is being forcefully removed, suggesting it may perceive such forceful actions as violent. In the second example, despite several interactive groups being present, the CAMs concentrate on a collision involving a woman and a group, which might have been a violent scenario in reality. These observations indicate that, while the model has effectively learned to pinpoint cues related to violent actions, it may sometimes overemphasize certain visual signals too.
@@ 4.5 Ablation Study
We perform a series of ablation studies to assess the efficacy of the components of our architecture on the combined dataset as it provides a more balanced class distribution and incorporates multiple sources and domains, which helps reduce dataset specific biases and focus on the model instead.
4.5.1 Ablation on Cropping Module and Residual Skip Connections
Model Cropping - $\times$ Cropping - $\checkmark$
Skip - $\times$ Skip - $\checkmark$ Skip - $\times$ Skip - $\checkmark$
VideoMamba - M (Spatial First) 92.90% 94.00% 93.44% 94.54%
VideoMamba - L (Spatial First) 92.46% 92.90% 93.11% 93.55%
VideoMamba - M (Temporal First) 91.8% 92.79% 92.24% 93.01%
Dual Branch (Ours) 94.64% 95.01% 95.19% 95.85%
Table 4: Performance comparison on accuracy with and without the cropping module and the residual skip connections among the blocks in each branch.
Tab. 4 compares model performance with and without the cropping module and the residual skip connections among the blocks in each branch. We observe that all the model variants benefit from the cropping mechanism, as it focuses the network on regions where people are most likely to appear, along with the residual skip connections. However, even though these mechanisms are also present in standard VideoMamba, including its Large variant with a comparable parameter size, the performance still lags behind our Dual Branch design. These results also indicate that relying on spatial or temporal scanning alone is insufficient, while our architecture leverages both more effectively to achieve superior results.
4.5.2 Ablation on Fusion Mechanism in Lateral Connections
@@
Concatenated LCs (Full Hidden State) 76.17
Concatenated LCs (CLS Token Only) 94.64
Additive LCs 93.44
Cross Attention based LCs 95.19
Gated LCs (Branch-2 $\rightarrow$ Branch-1) 93.22
Gated LCs (Branch-1 $\rightarrow$ Branch-2) 95.85
Table 5: Performance comparison of different fusion mechanisms.
In Tab. 5, we compare four fusion mechanisms for lateral connections: concatenation-based, additive-based, cross-attention based and a gated approach. Since concatenating full hidden states resulted in a severe drop in the accuracy, apparently because the token spaces are misaligned across scan orders, causing redundant/scale-mismatched features, all the other experiments were conducted based on fusing only the CLS Tokens from each branch as it seemed, selective fusion via a bottleneck avoids the redundancy and misalignment that may have arisen when fusing full hidden states. Gated Lateral Connections from Branch-1 (spatial-first scanning) to Branch-2 (temporal-first scanning) outperformed all other methods, because spatial cues might have provided strong priors that help the temporal branch to focus on meaningful motion patterns.
4.5.3 Ablation on Continuous Fusion
Lateral-Connection (LC) Config. Accuracy (%)
LCs alternatively (even layers) 95.63
LCs alternatively (odd layers) 95.63
One LC only at beginning 94.00
One LC only at end 94.43
One LC at middle 94.32
Two LCs at beginning and end 95.30
Continuous LCs 95.85
Table 6: Performance comparison for different configurations of lateral connections.
In Tab. 6, we investigate the optimal configuration for lateral connections between the two branches. Various strategies including applying lateral connections at different layers (e.g., early, middle, late) or alternately across even/odd layers are investigated. While all configurations exceed 94% accuracy, continuous lateral connections (i.e., applying them at every block) yield the highest accuracy of 95.85%. This finding highlights the importance of gradually fusing information between the two branches continuously throughout the network rather than restricting fusion to early or late stages.
@@ 5 Conclusion
In this paper, we present a novel Dual Branch VideoMamba architecture with Gated Class Token Fusion (GCTF) for violence detection in videos. Our method integrates an efficient state-space model with a dual-stream design aimed at separately capturing fine-grained spatial features and global temporal dynamics. By continuously fusing the class tokens from both branches using a learnable gating mechanism, our approach effectively combines complementary cues, yielding state-of-the-art performance. Extensive experiments on the combined benchmark dataset and the DVD dataset, demonstrate that our model not only achieves state-of-the-art accuracies but also shows a significant reductions in model parameters and computational cost in FLOPS compared to previous state-of-the-art methods such as CUE-Net [35] and VideoMamba [26] variants. This work highlights the promise of state-space models for scalable and near real-time video violence detection and paves the way for future research, including the integration of multi-modal data and further refinement of the fusion strategies.
@@
Acknowledgment. This work has been funded through an EPSRC DTP studentship at Queen Mary University of London. This paper utilized Queen Mary’s Andrena HPC facility, supported by QMUL Research-IT Services [21].
Supplementary Material
5.1 Cropping Module - Supplement
The cropping mechanism in the cropping module extracts the region that encompasses all detected people in each frame, to focus specifically on the violent actions taking place, based on the observation that violent incidents typically involve multiple humans. As seen in Fig. 5, by computing the maximum bounding area around all detected people, the network is guided to the most informative spatial areas while retaining the original frames if no individuals are detected. Temporal cropping is also not applied here to prevent further potential information loss from missing undetected individuals. Following the practice from [35], YOLO (You Only Look Once) V8 algorithm [19] which classifies objects in a single pass using a CNN-based architecture was used to detect people in this cropping module.
Figure 5: The cropping module which includes YOLO V8 algorithm to detect the people present and crop the maximum bounding region.
5.2 Individual DataSet Descriptions - Supplement
5.2.1 Real World Fighting (RWF-2000) Dataset
The Real World Fighting (RWF-2000) dataset [9] was introduced in 2020 and is a comprehensive dataset that contains real world fighting scenarios sourced purely through surveillance footage. This dataset contains 2,000 trimmed video clips and each video is trimmed into 5 seconds where the fighting occurs. The dataset is balanced, with a 80%/20% train-test split which has been thoroughly checked for data leakage between the splits.
5.2.2 Real Life Violence Situations (RLVS) Dataset
The Real Life Violence Situations (RLVS) dataset [36] consists of 2000 video clips with 1000 violent and another 1000 non-violent videos collected from YouTube. These contain many real street fight situations in several environments with an average length of 5s from different sources such as surveillance cameras, movies, video recordings, etc. Similar to the RWF-2000 dataset, a 80%/20% train-test split has been created for this dataset as well.
5.2.3 Vision-based Fight Detection From Surveillance Cameras (SURV) Dataset
SURV [2] is a publicly available dataset created specifically for real-world fight detection from surveillance footage. The videos were collected from YouTube surveillance camera sources and contain fight and non-fight scenes captured in unconstrained environments such as streets, public areas, and institutions. Each video is trimmed into short clips ranging from 2s - 3s and a 80%/20% train-test split has been created for this dataset as well. Unlike staged datasets (e.g., movies, hockey), SURV reflects authentic surveillance conditions, including challenges like occlusion, varying viewpoints, and low resolution
5.2.4 VioPeru Dataset
In [16], the researchers have created a new balanced dataset called VioPeru, which consists of 280 videos collected from real video surveillance camera records containing challenging violent incidents involving two or more people. The videos have been collected from the citizen security offices of different municipalities in Peru. The videos have been trimmed to 5s just to include the violent incident. Similar to the above datasets, a 80%/20% train-test split has also been created.
5.2.5 DVD Dataset
In our DVD pipeline we did not use fixed windows; instead, we directly segmented from the frame-level annotations by scanning each video’s label stream (aligned via FPS to timestamps) for maximal contiguous runs of the same class and turning those runs into clips: runs labeled 0 were recorded as violent clips and runs labeled $1$ as non-violent clips, while any frames labeled 1-1 (irrelevant/uncertain) were treated as “ignore.” Whenever a 1-1 label occurred, the corresponding portion was discarded rather than assigned to either class, and any runs bordering a 1-1 region were truncated at the boundary so no ambiguous frames leaked into the clips. For each retained run we stored start and end times, duration, FPS, width/height, the source file name, and a stable relative path in a manifest, ensuring clips never overlap across classes and that every exported segment is a faithful, contiguous slice of unambiguous labels derived directly from the annotations. Media extraction was performed only after the manifest was finalized (late binding), snapping cut points to sensible decoding boundaries; no gap-merging or window resampling was applied, and no re-labeling heuristics were used—clip identities come solely from contiguous $0/1$ runs with all 1-1 spans excluded.
[…]
5.2.6 Data Leakage Findings after Amalgamation:
SURV vs. RWF-2000 and RLVS: We find no similar videos between the SURV test set and either RWF-2000 or RLVS training sets despite SURV dataset being collected from publicly published surveillance videos in social media.
VioPeru vs. RWF-2000 and RLVS: We find no similar videos between the VioPeru test set and either RWF-2000 or RLVS training sets. This is not unexpected, as VioPeru is a newly collected dataset from Peruvian municipalities and contains unique CCTV footage which has not been released on YouTube or any other social media platform.
RWF-2000 and RLVS: We discover one instance where a video in the testing set of RLVS is identical to a video in the training set of RWF-2000. To prevent leakage, we remove the duplicate entry from the RLVS testing set.
6 Implementation Details - Supplement
Our Dual Branch VideoMamba architecture was implemented in PyTorch using the AdamW optimizer [29] with a cosine learning rate schedule starting with a learning rate of 1e-4 and Cross-Entropy Loss, taking insights from training recipes of the original VideoMamba architecture. All models were trained for 55 epochs with 5 warm-up epochs where the best validation model was saved.
7 Dual Branch Video Mamba Performance on separate datasets of RWF-2000, RLVS and VioPeru
When trained and tested separately, our architecture outperforms the reported state-of-the-art results in literature in classification accuracy for RWF-2000, RLVS and SURV datasets, by achieving accuracies of 94.50%, 99.75% and 96.67% respectively, setting a new state-of-the-art. For the VioPeru dataset, our Dual Branch architecture is able to reach the already reported state-of-the-art accuracy of 89.23%.
Dataset Reported Best Accuracy (%) Our Model Accuracy (%)
RWF-2000 [9] 94.36 [7] 94.50
VioPeru [16] 89.23 [16] 89.23
RLVS [36] 99.50 [35] 99.75
SURV [2] 95.62 [40] 96.67
Table 7: Comparison of earlier best accuracies with Dual Branch VideoMamba on three datasets.
Method Model Type Accuracy (%)
CNN- LSTM [36] VGG16+LSTM 88.20
Temporal Fusion CNN +LSTM [10] CNN+LSTM 91.02
DeVTr [1] ViViT 96.25
ACTION- VST [23] CNN + ViViT 98.69
CUE-Net Enhanced UniformerV2 99.50
Video- Mamba SSM 99.50
Dual Branch (Video- Mamba) SSM 99.75
Table 8: Results comparison for the RLVS Dataset.
Method Model Type Accuracy (%)
ConvLSTM [9] CNN+LSTM 77.00
X3D [23] 3DCNN 84.75
I3D [15] 3DCNN 83.40
Flow Gated Network [9] Two Stream Graph CNN 87.25
SPIL [38] Graph CNN 89.30
Structured Keypoint Pooling [15] CNN 93.40
Video Swin Transfor- mer [28] ViViT 91.25
ACTION- VST [23] CNN + ViViT 93.59
CUE-Net Enhanced UniformerV2 94.00
Video- Mamba SSM 92.75
Multi- Head Att & LSTM [7] LSTM + ViViT 94.36
Dual Branch (Video- Mamba) SSM 94.50
Table 9: Results comparison for the RWF-2000 Dataset.
Method Model Type Accuracy (%)
Sep Conv [16] LSTM- CNN + LSTM 73.21
Advanced Sep Conv [16] LSTM CNN + LSTM 89.29
Video- Mamba SSM 85.71
Dual Branch (Video- Mamba) SSM 89.29
Table 10: Results comparison for the VioPeru Dataset.
Method Model Type Accuracy (%)
Temporal Spatial [45] Attn. Maps CNN + Attn 91.80
RTFM  [40] MIL based 95.62
Video- Mamba SSM 95.00
Dual Branch (Video- Mamba) SSM 96.67
Table 11: Results comparison for the SURV Dataset.
7.1 Ablation Study - Supplement
7.1.1 Ablation on Number of Frames for each Branch
Tab. 12 examines the impact of varying the number of frames inputted to each branch. This ablation study shows that providing each branch with the same frame count appears beneficial when it comes to our Dual Branch VideoMamba architecture, likely due to more balanced representation learning and consistent temporal context across branches.
Branch-1 Branch-2 Accuracy (%)
32 64 94.75
64 32 95.08
32 32 95.74
64 64 95.85
Table 12: Comparison of model performance with different number of frames inputted into 2 branches evaluated against the combined dataset.
7.1.2 Ablation on Fusion Mechanism in Final Fusion Block
[…]
Fusion Mechanism Accuracy (%)
Addition 94.97
Cross Attention 95.74
Gated LCs (Branch-2 $\rightarrow$ Branch-1) 93.77
Gated LCs (Branch-1 $\rightarrow$ Branch-2) 95.30
Concatenation 95.85
Table 13: Performance comparison of different fusion mechanisms in the final fusion block for the combined dataset.